OpenStudy (hockeychick23):

You choose to participate in a raffle. Raffle tickets sell for $5 each and there are a total of 10,000 tickets. The prizes in this raffle are: 1st prize: $20,000 college scholarship (1 winner) 2nd prize: $10,000 college scholarship (2 winners) 3rd prize: $1,000 college scholarship (5 winners) 4th prize: $100 college scholarship (50 winners) You purchase one $5 raffle ticket. How much of your college tuition can you expect to be paid for by your raffle tickets purchase?

OpenStudy (hockeychick23):

@ganeshie8

OpenStudy (hockeychick23):

@satellite73

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

there is a probability way to do this but there is another easy way too

OpenStudy (misty1212):

buy ALL the tickets how much do you spend?

OpenStudy (hockeychick23):

all the tickets would cost $50,000

OpenStudy (misty1212):

right how much money do you get back? in total?

OpenStudy (misty1212):

@iambatman yea how did you guess?

OpenStudy (hockeychick23):

$50,000 in scholarships?

OpenStudy (misty1212):

hmm no lets add up all the prize money

OpenStudy (misty1212):

\[20,000+20,000+3000+400=43,400\]

OpenStudy (misty1212):

you spend \(50,000\) you get back \(43,400\) if my arithmetic is correct, for a net loss of \[43400-50000=6600\]

OpenStudy (misty1212):

average that over the \(10,000\) tickets and get \[-\frac{6600}{10000}\]

OpenStudy (misty1212):

or \(-.66\) i.e. you expect to lose \(.66\) per ticket

OpenStudy (hockeychick23):

so they only lose 66 cents? The answer choices I'm given are: A. -$10.20 B. $5.00 C. $4.56 D.-$3.72 E. $100

OpenStudy (misty1212):

maybe my arithmetic was bad let me start again

OpenStudy (misty1212):

i see, it was the wording of the question you pay $5, you lose \(.66\) so they are saying it is \(\$4.54\)

OpenStudy (misty1212):

that is the net loss of \(.66\)

OpenStudy (hockeychick23):

But how do i get from 66 cents to a much higher number? because the answers are much bigger numbers?

OpenStudy (misty1212):

not sure what you mean you expect to lose 66 cents per ticket, making your $5 ticket worth \(5.00-.66=\$4.54\)

OpenStudy (hockeychick23):

hmm im not sure why its not one of the answers then :(